Circular No. 7844

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bg IN MCG +02-38-31 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2002bg (cf. IAUC 7836), obtained by P. Berlind on Mar. 6.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova two weeks after maximum.

SUPERNOVA 2002bh IN UGC 5286 S. Benetti, G. Altavilla, A. Pastorello, M. Turatto, and S. Desidera,

Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A noisy spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2002bh (cf. IAUC 7837), obtained on Mar. 5.92 UT with the Asiago 1.82-m telescope (+ AFOSC), shows that it is a type-II supernova, a few days after maximum. The spectrum is dominated by a very broad (FWHM ∼ 14000 km/s) Hα emission.”

Matheson et al. also communicate that a spectrum of SN 2002bh, obtained by Berlind on Mar. 6.24 UT as above, shows it to be a type-II supernova near maximum. The spectrum consists of a relatively blue continuum with P-Cyg lines of hydrogen and helium. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5198 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 10 400 km/s.

SUPERNOVA 2002bj IN NGC 1821 K. Kinugasa and H. Kawakita, Gunma Astronomical Observatory

(GAO); and H. Yamaoka and H. Naito, Kyushu University, report that a low-resolution spectrum of SN 2002bj (IAUC 7839) was obtained with the GAO 0.65-m telescope (+ GSC) on Mar. 4.49 UT, showing a blue continuum with a wide and shallow absorption feature at 620 nm. The spectrum closely resembles that of type-II SN 2001hg before maximum (IAUC 7770).

Matheson et al. add that a spectrum of SN 2002bj, obtained by Berlind as above on Mar. 6.11 UT, shows it to be a type-IIn supernova. The spectrum consists of a very blue continuum with P-Cyg lines of hydrogen and helium, similar to that of SN 1998S, 25 days after discovery (see Leonard et al. 2000, Ap.J. 536, 239), but without the narrow P-Cyg features.

2002 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green